An airbag is filled with 18L nitrogen gas at 1 atm and 27 degrees celsius, what mass of NaN2 must be decomposed?
Answer #1

Given:

P = 1.0 atm

V = 18.0 L

T = 27.0 oC

= (27.0+273) K

= 300 K

find number of moles using:

P * V = n*R*T

1 atm * 18 L = n * 0.08206 atm.L/mol.K * 300 K

n = 0.7312 mol

This is number of mol of N2 produced

The reaction taking place is:

NaN2 —> Na   +   N2

from reaction,

mol of NaN2 required = mol of N2 formed

= 0.7312 mol

Molar mass of NaN2,

MM = 1*MM(Na) + 2*MM(N)

= 1*22.99 + 2*14.01

= 51.01 g/mol

use:

mass of NaN2,

m = number of mol * molar mass

= 0.7312 mol * 51.01 g/mol

= 37.3 g

Answer: 37.3 g

Answer: 37.3 g